Sam is 30 years older than Bob in 10 years from now Sam will be twice as old as Bob will be then find the present ages
Answer by stanbon(75887)

Sam is 30 years older than Bob.
10 years from now Sam will be twice as old as Bob.
Find the present ages
----
Equations:
s = b + 30
s+10 = 2(b+10)
-----
Substitute for "s" and solve for "b"::
b+30+10 = 2b+20
b = 20 yrs (Bob's age now)
------
Solve for "s":
s = b + 30
s = 20+30 = 50 yrs (Sam's age now)
----------
Cheers,
Stan H.